There is 60% chance that a final year student would throw a party before leaving school, taking every 50 student from a total of 150. Find the mean value and the variance.
1
Expert's answer
2021-07-26T09:21:34-0400

Solution:

Given, "p=60\%=0.6, n=50"

"Mean=\mu=np=0.6\times50=30
\\Variance=\sigma^2=np(1-p)=0.6(50)(0.4)=12"